Use the Newton’s method to approximate the real solution of xcube-2x-2=0 in the interval [1,2]


1
Expert's answer
2021-06-01T18:16:58-0400

x3 - 2x - 2 = 0 x "\in" [ 1, 2 ]


Let f(x) = x3 - 2x - 2


f(1) = (1)3 - 2(1) - 2

f(1) = -3



f(2) = (2)3 - 2(2) - 2

f(2) = 2


Hence, a real solution of f(x) exists between 1 and 2.


f ' (x) = 3x2 - 2


Let x0 = 1.5



xn+1 = xn - "\dfrac{f(x\scriptscriptstyle n)}{f'(x\scriptscriptstyle n)}"


"\implies" n= 0


x1 = x0 - "\dfrac{f(x\scriptscriptstyle 0)}{f'(x\scriptscriptstyle 0)}"



x1 = 1.5 - "\dfrac{f(1.5)}{f'(1.5)}"


f(1.5) = (1.5)3 - 2(1.5) - 2 = -1.625


f ' (x) = 3(1.5)2 - 2 = 4.75



x1 = 1.5 + "\dfrac{1.625}{4.75}"


x1 = 1.842





"\implies" n =1



x2 = x1 - "\dfrac{f(x\scriptscriptstyle 1)}{f'(x\scriptscriptstyle 1)}"



x2 = 1.842 - "\dfrac{f(1.842)}{f'(1.842)}"


f(1.842) = (1.842)3 - 2(1.842) - 2 = 0.566


f ' (x) = 3(1.842)2 - 2 = 8.178



x2 = 1.842 - "\dfrac{0.566}{8.178}"


x2 = 1.772








"\implies" n =2



x3 = x2 - "\dfrac{f(x\scriptscriptstyle 2)}{f'(x\scriptscriptstyle 2)}"



x3 = 1.772 - "\dfrac{f(1.772)}{f'(1.772)}"


f(1.772) = (1.772)3 - 2(1.772) - 2 = 0.0201


f ' (x) = 3(1.772)2 - 2 = 7.420



x3 = 1.772 - "\dfrac{0.0201}{7.420}"


x3 = 1.769






"\implies" n = 3



x4 = x3 - "\dfrac{f(x\scriptscriptstyle 3)}{f'(x\scriptscriptstyle 3)}"



x4 = 1.769 - "\dfrac{f(1.769)}{f'(1.769)}"


f(1.769) = (1.769)3 - 2(1.769) - 2 = -0.00216


f ' (x) = 3(1.769)2 - 2 = 7.388



x4 = 1.769 - "\dfrac{-0.00216}{7.388}"


x4 = 1.769





Now we see that x4 = x3 . Hence, the root of the given equation is 1.769 correct up to three decimal places.